OpenStudy (spareb665):

Write the zeros of the polynomial given below. f(x)= 1/3x^3 + 2x^2 +x - 10/3 = 1/3 (x+5)(x-1)(x+2)

OpenStudy (campbell_st):

since its been factorised all you need to do for the zeros is solve x + 5 = 0 x -1 =0 x + 2 = 0

OpenStudy (campbell_st):

well the right hand side of the equation has been factorised so that f(x) = 1/3(x + 5)(x -1)(x + 2) if any of the binomials (x+5), (x -1) or (x + 2) is equal to zero then then f(x) will be equal to zero. this is because you are multiplying 3 binomials if any binomial is zero then the produce is zero. So you are finding the values of x where the function cuts the x- axis... this is where y = 0... hence the term... find the zeros. so all you do is solve x + 5 = 0 so x = -5 for it to be true you need to solve x - 1 = 0 and x + 2 = 0 hope this helps...

OpenStudy (anonymous):

http://www.youtube.com/watch?v=cGaps65Y6lo